# Rethinking Approximate Gaussian Inference in Classification

### v1

@@ 2.3 Making the Approximations Exact
Let $\varphi=\rho$. Suppose that the logit means and variances $(\mu_{c},\sigma^{2}_{c})$ lie in some compact set $\mathcal{K}\subset\mathbb{R}\times\mathbb{R}_{>0}$ for each class $c$. Then, assuming the constraint Equation 18 is perfectly satisfied, the KL divergence between the true and approximate predictive Equation 17 is bounded by some constant $M(\mathcal{K})$ only depending on $\mathcal{K}$, not on $C$.
Section C.2 provides a more precise statement, a formula for $M(\mathcal{K})$, and a proof. Importantly, this result states that the approximation error does not increase with the number of classes.
Finally, observe that, when enforcing Equation 18, we obtain exact predictives with Equation 13, which only uses the diagonal of $\mathrm{f}(x)$’s covariance. Enforcing this constraint thus allows one to ignore non-diagonal terms in the logit covariances. In Section 5, we discuss how this can be leveraged for computational gains.
@@ 3.1 Dirichlet Matching
The formulation of Equation 11 does not yet allow this. However, we will now see that we can obtain good Dirichlet approximations by constructing a tractable approximation of Equation 11 of the form
@@
where $\mathcal{D}(\Delta^{C-1})$ is the space of Dirichlet distributions on $\Delta^{C-1}$. The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest can be computed analytically (c.f. Appendix G). The map $\mathrm{a}$ is constructed by moment matching: we match the first two moments of the Dirichlet distribution to the moments of the Gaussian pushforward $\bm{\varphi}_{*}\mathrm{f}(x)$.
Conveniently, just as the first moments described in Section 2, we can obtain second moments for $\exp$
@@
and for $\Phi$ (Owen (1980, Eq. 20,010.4))
$\int_{\mathbb{R}}\Phi(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (21)
$\quad=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)-2T\left(\frac{\mu}{\sqrt{1+\sigma^{2}}},\frac{1}{\sqrt{1+2\sigma^{2}}}\right),$
where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau (2017, Equation 23))
$\int_{\mathbb{R}}\rho(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)$ (22)
$-\frac{\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\left(1-\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\right)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}.$
@@
$\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{P}^{2}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]$ (23)
$\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}^{2}]}{\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]^{2}}.$
As in Equation 13, the approximation in Equation 23 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable, as under the constraint of Equation 18.
If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Equations 13 and 23, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns. Thus, we instead use a classical Dirichlet method of moments (Minka, 2000, Eqs. 19 & 23), which reduces these equations, as we will now describe.
@@
for any $1\leq c\leq C$. So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Equation 24. Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Equation 24 an expression for the Dirichlet parameters
@@
Hence, using Equations 13 and 23, we obtain the computationally feasible Dirichlet parameters
@@
for $\mathrm{h}(x)$, where $S=\max(\sum_{c=1}^{C}\mathbb{E}[Q_{c}],1)$. Taking the maximum with $1$ in the expression for $S$ is an additional approximation to ensure $\bm{\gamma}\geq\bm{0}$, and is justified when taking $s(x)=1$ for all $x$ in Equation 18. Note that performing such moment matching to obtain the Dirichlet parameters in Equation 26, the mean of this Dirichlet matches precisely the approximate predictive Equation 13. That is, the approximate predictive of the exact distributional model (Equation 11) matches the exact predictive of the approximate distributional model (Equation 19).
@@ 3.2 Intermediate Beta Matching
When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Equation 19) further:
@@
where $\mathcal{B}((0,1))$ is the space of Beta distributions on $(0,1)$. Like the output Dirichlets, we obtain the intermediate Beta distributions by moment matching (Appendix D). Figure 2 indicates that such Beta approximations tend to be of high quality. In Section D.1, we give an information-geometric motivation for this, which shows that the choice of activation $\Phi$ is ideal for approximating the Gaussian pushforwards with Beta distributions.
The quality of such Beta approximations is encouraging for that of the output Dirichlet approximations: observe that taking $s(x)=1$ for all $x$ in Equation 18, $\bm{n}$ becomes the identity (as it normalises by 1), and thus all $\mathrm{n}$ does is combining the $C$ Beta approximations into one Dirichlet.
4 Learning the Constraint
As discussed in Sections 2.3 and 3, the quality of our approximations relies on the enforcement of the constraint $\sum_{c=1}^{C}Q_{c}=s(x)$ for some deterministic function $s\colon\mathcal{X}\to\mathbb{R}$ (Equation 18). In this work, we simply choose $s(x)=1$ for all $x$. We thus make use of the degree of freedom in the logits to obtain exact predictives in Equation 13. In Appendix E, we make the connection with the energy-based model view of classifiers (Grathwohl et al., 2019), which instead makes use of this degree of freedom to turn discriminative models into generative models. Future work could combine both frameworks by learning $s(x)\propto p_{\operatorname{gen}}(x)$.
4.1 Explicit Constraint
Up to this point, we have treated the model $\mathrm{f}$ as a black box. However, to discuss how to impose Equation 18, we need to think about the specifics of how such models are trained and how Gaussian distributions are then obtained.
@@ 4.1 Explicit Constraint
In linearised Laplace approximations and SNGP, a neural network $\bm{f}\colon\mathcal{X}\to\mathbb{R}^{C}$ is trained for the mean of $\mathrm{f}$, while the covariance is obtained post-hoc. Therefore, to enforce Equation 18 on the training data distribution, we can train for $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$ for all $1\leq n\leq N$.
This can be achieved, for example, by adding an appropriate regulariser to the loss. Namely,
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ (28)
=~$\quad=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))-1\right)^{2}$
for some tuned regularisation constant $\lambda$, where $\tilde{\mathcal{L}}$ is the original unregularised loss.
In contrast to the previously mentioned methods, HET directly trains not only for the mean $\bm{f}$ but also for the covariance matrix $\bm{\Sigma}$ of $\mathrm{f}$ through the predictives. Thus, in this case, we can regularise with the means of the Gaussian pushforwards through $\varphi$:
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ =~$=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})$ (29)
$+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}]-1\right)^{2}.$
In Section 6, we will see that training with such a learning objective does not always outperform vanilla CE training. For $\varphi=\exp$, for instance, we observe that the regulariser can take on extreme values, dominating the data term, slowing down training.
We shall now see that, when $\varphi$ has range $(0,1)$, we can enforce the constraint with the learning objective in a more natural, implicit way.
4.2 Implicit Constraint
In the case of $\varphi=\Phi\text{ or }\rho$, $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$ can be achieved more advantageously: since the range of $\Phi$ and $\rho$ is $(0,1)$, under the condition $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$, we can view $\varphi(f_{c}(x_{n}))$ as estimates of $p(c\mid x_{n})$. This motivates the use of class-wise cross-entropy losses (class-wise CE).
For linearised Laplace approximations and SNGP, which solely train a mean $\bm{f}$, the class-wise CE is given by
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big{(}\delta_{c_{n},c}\log(\varphi(f_{c}(x)))$ (30)
$\quad+(1-\delta_{c_{n},c})\log(1-\varphi(f_{c}(x))\big{)}.$
The labels being one-hot vectors and the binary cross-entropy loss being a (negative) strictly proper scoring rule (c.f. Appendix I) together ensure that Equation 30 enforces the soft constraint $\sum_{c=1}^{C}\varphi\circ\mu_{c}=1$ on the training data, and hence in the training data distribution (Figure 3).
We note that the loss Equation 30 is only used for training and not for inference or evaluation. Indeed, methods such as Laplace approximations and SNGP – which rely on negative log-likelihood losses to obtain an approximate Gaussian distribution – can then use the actual log-likelihood loss of the model, that is, the cross-entropy loss of $\bm{n}\circ\bm{\varphi}\circ\bm{f}$.
Similarly, for HET, we use the class-wise CE with the pushforward means
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big{(}\delta_{c_{n},c}\log(\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}])$ (31)
$+(1-\delta_{c_{n},c})\log(1-\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}])\big{)}.$
Enforcing the constraint in this implicit way is powerful since, in contrast to the regularised CE objective (Equation 28), the class-wise CE (Equation 30) fully aligns with the task. In Section 6 we shall see that training on class-wise CE generally outperforms softmax CE training.
Figure 3: Evolution of the constraint on validation data for normCDF models trained on ImageNet.
@@ 5 Computational Gains
The runtime cost of the MC predictive approximation is $\mathcal{O}(S\cdot C+C^{3})$, and the memory cost is $\mathcal{O}(S\cdot C+C^{2})$, where $S$ is the number of samples. The $C^{2}$ term in memory comes from the logit covariance matrix; the $C^{3}$ term in runtime comes from the need to Cholesky-decompose such covariance matrix before sampling. For example, on a ResNet-50, sampling one thousand logit vectors from the logit-space Gaussian means an approximately 7% overhead on the forward pass. In embedded systems with much smaller models, this overhead is even larger. In contrast, the runtime and memory cost of calculating a predictive through our framework is $\mathcal{O}(C)$, which is negligible in comparison.
Indeed, as noted at the end of Section 2.3, the constraint (Equation 18) ensures that only the mean and the diagonal of the logit covariance are needed in order to calculate the predictive exactly. That is, the logit quantities required to compute a predictive scale as $\mathcal{O}(C)$ instead of $\mathcal{O}(C^{2})$. This provides further opportunities for computational savings in the methods used to obtain the Gaussian distributions.
In HET, for example, the logit covariance is trained with both a diagonal term and a low-rank term. Our framework thus allows one to drop the low-rank term without losing any expressivity in the model.
In SNGP, the off-diagonal terms of the logit covariance are ignored (Liu et al., 2023, p. 47). For unconstrained models, this incurs an approximation error as the inter-class covariance structure can encode vital information for inference. When Equation 18 is satisfied, however, we are guaranteed not to lose expressivity for the predictive computation by discarding the off-diagonal covariance terms, theoretically backing the approximation in SNGP.
@@ 6 Experiments
  • Do we have to sacrifice performance for sample-free predictives?
  • What are the effects of changing the learning objective?
To answer the first question, we consider fixed (method, activation) pairs and verify that our methods perform on par with the Monte Carlo sampled predictives in Section 6.1.
We train four classes of models: Softmax ($\varphi=\exp$, trained with vanilla CE), Exp ($\varphi=\exp$, trained with the explicit constraint (Section 4.1)), NormCDF ($\varphi=\Phi$, trained with the implicit constraint, (Section 4.2)) and Sigmoid ($\varphi=\rho$, trained with the implicit constraint, (Section 4.2)).
We consider Heteroscedastic Classifiers (HET) (Collier et al., 2021), Spectral-Normalized Gaussian Processes (SNGP) (Liu et al., 2020), and last-layer Laplace approximation methods (Daxberger et al., 2021) as backbones. The resulting 18 (method, activation, predictive) triplets are evaluated on ImageNet-1k (Deng et al., 2009) and CIFAR-10 (Krizhevsky & Hinton, 2009) on five metrics aligning with practical needs from uncertainty estimates (Mucsányi et al., 2023). Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples. For ImageNet, we treat ImageNet-C (Hendrycks & Dietterich, 2019) samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
For the second question, we evaluate our analytic predictives and moment-matched Dirichlets against softmax models with approximate inference tools – Laplace Bridge (Hobbhahn et al., 2022) and Mean Field (Lu et al., 2021) approximations, as well as MC sampling, in Section 6.2.
To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones. For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al. (2024). On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. We search for ideal hyperparameters and checkpoints with a ten-step Bayesian Optimization scheme (Shahriari et al., 2015) in Weights & Biases (Biewald, 2020).
The main paper focuses on ImageNet results to highlight the scalability of our framework and only shows proper scoring results for CIFAR-10. For a complete overview of CIFAR-10 results, refer to Appendix K.
In all plots of the main paper and the Appendix, the error bars are calculated over five independent runs with different seeds and show the minimum and maximum performance.
@@
MC 10 0.0112 0.0005
Analytic 0.0114 0.0007
MC 100 0.0115 0.0006
MC 1000 0.0118 0.0007

### v2

@@ 2.3 Theoretical Analysis of the Approximation’s Quality
Suppose the means and variances $(\mu_{c},\sigma_{c}^{2})$ lie in some compact set $\mathcal{K}\subset\mathbb{R}\times[0,\infty)$ for each class $c$. Then
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})=M(\mathcal{K})+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (19)
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, for some constant $M(\mathcal{K})\in[0,\infty]$ which only depends on $\mathcal{K}$, not on $C$, such that $M(\mathcal{K})<\infty$ when $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$, and $M(\mathcal{K})=0$ when $q=\hat{q}$ for all $(\mu,\sigma^{2})\in\mathcal{K}$.
Section C.2 provides a formula for $M(\mathcal{K})$ and a proof. We also conduct an empirical investigation of whether the assumptions made in the theorem hold in practice, namely whether the logit means and variances lie in a compact set, and whether $M(\mathcal{K})<\infty$.
The $O$ term in Eq. 19 is due to the splitting of the expectation into the numerator and denominator in the derivation Eq. 10, while $M(\mathcal{K})$ results from the error due to the probit approximation. Importantly, Theorem 2.1 states that the latter does not increase with the number of classes.
@@ 3.1 Dirichlet Matching
The formulation of Eq. 11 does not yet allow this. However, we will now see that one can obtain good Dirichlet approximations by constructing a tractable approximation of Eq. 11 of the form
@@
where $\mathcal{D}(\Delta^{C-1})$ is the space of Dirichlet distributions on $\Delta^{C-1}$. The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest are closed-form (see Appendix F). The map $\mathrm{a}$ is constructed by moment matching: we match the first two moments of the Dirichlet distribution to the moments of the Gaussian pushforward $\bm{\varphi}_{*}\mathrm{f}(x)$.
Conveniently, like the first moments described in Section 2, we can obtain second moments for $\exp$
@@
and for $\Phi$ (Owen [30, Eq. 20,010.4])
$\int_{\mathbb{R}}\Phi(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)-2T\left(\frac{\mu}{\sqrt{1+\sigma^{2}}},\frac{1}{\sqrt{1+2\sigma^{2}}}\right),$ (22)
where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau [7, Equation 23])
$\int_{\mathbb{R}}\rho(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)-\frac{\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\left(1-\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\right)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}.$ (23)
@@
$\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{P}^{2}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}^{2}]}{\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]^{2}}.$ (24)
As in Eq. 12, the approximation in Eq. 24 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable.
If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs. 12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns. Thus, we instead use a classical Dirichlet method of moments [25, Eqs. 19 & 23], which reduces these equations, as we will now describe.
@@
for any $1\leq c\leq C$. So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Eq. 25. Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Eq. 25 an expression for the Dirichlet parameters
@@
Hence, using Eqs. 12 and 24, we obtain the computationally feasible Dirichlet parameters
@@
for $\mathrm{h}(x)$, where $S=\max(\sum_{c=1}^{C}\mathbb{E}[Q_{c}],1)$. Taking the maximum with $1$ in the expression for $S$ is an additional approximation to ensure $\bm{\gamma}>\bm{0}$. Note that by performing such moment matching to obtain the Dirichlet parameters in Eq. 27, the mean of this Dirichlet matches precisely the approximate predictive Eq. 12. That is, the approximate predictive of the exact distributional model (Eq. 11) matches the exact predictive of the approximate distributional model (Eq. 20).
@@ 3.2 Intermediate Beta Matching
When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Eq. 20) further:
@@
where $\mathcal{B}((0,1))$ is the space of Beta distributions on $(0,1)$. Like the output Dirichlets, we obtain the intermediate Beta distributions by moment matching (Appendix D). Figure 2 indicates that such Beta approximations tend to be of high quality. In Section D.1, we give an information-geometric motivation for this, which shows that the choice of activation $\Phi$ is ideal for approximating the Gaussian pushforwards with Beta distributions. The quality of such Beta approximations is encouraging for that of the output Dirichlet approximations. Analysing this theoretically presents an interesting direction for future research.
4 Training
With the choices $\varphi=\Phi$ or $\rho$, we could still train the underlying model with a cross-entropy (CE) loss, replacing the softmax by the activation $\bm{a}=\bm{n}\circ\bm{\Phi}$ or $\bm{n}\circ\bm{\rho}$. However, we can instead leverage the fact that $\Phi$ and $\rho$ have values in $(0,1)$ to train using the binary cross-entropy (BCE) loss.
The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix H). Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [37, 19]. The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section 6).
@@ 4 Training
For example, in linearised Laplace and SNGP, a neural network $\bm{f}\colon\mathcal{X}\to\mathbb{R}^{C}$ is trained for the mean of $\mathrm{f}$, while the covariance is obtained post-hoc. Thus, for these methods, we can train the mean function $\bm{f}$ with the BCE loss
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big{(}\delta_{c_{n},c}\log(\varphi(f_{c}(x)))+(1-\delta_{c_{n},c})\log(1-\varphi(f_{c}(x))\big{)}.$ (29)
@@ 5 Computational Gains
The runtime cost of the MC predictive approximation is $\mathcal{O}(S\cdot C+C^{3})$, and the memory cost is $\mathcal{O}(S\cdot C+C^{2})$, where $S$ is the number of samples. The $C^{2}$ term in memory comes from the logit covariance matrix; the $C^{3}$ term in runtime comes from the need to Cholesky-decompose such a covariance matrix before sampling. For example, on a ResNet-50, sampling one thousand logit vectors from the logit-space Gaussian means an approximately 7% overhead on the forward pass. In embedded systems with much smaller models, this overhead is even larger. In contrast, the runtime and memory cost of calculating our predictives is $\mathcal{O}(C)$, which is negligible in comparison.
Indeed, the approximation Eq. 12 uses only the mean and the diagonal of the logit covariance to calculate the predictive. That is, the logit quantities required to compute a predictive scale as $\mathcal{O}(C)$ instead of $\mathcal{O}(C^{2})$. This can provide further opportunities for computational savings in the methods used to obtain the Gaussian distributions.
In HET, for example, the logit covariance is trained with both a diagonal term and a low-rank term. Our framework thus allows one to drop the low-rank term (Section I.2).
@@ 6 Experiments
  1. (i)
    Do we have to sacrifice performance for sample-free predictives? (Section 6.1)
  2. (ii)
    What are the effects of changing the learning objective? (Section 6.2)
To answer the first question, we consider fixed (method, activation) pairs and verify that our methods perform on par with the Monte Carlo sampled predictives. In the main paper, we report results on Softmax ($\varphi=\exp$, trained with vanilla CE), NormCDF ($\varphi=\Phi$) and Sigmoid ($\varphi=\rho$) models. In Section K.1, we investigate the performance of Softmax models with closed-form predictives.
@@
MC 10 0.0112 $\pm$ 0.0005
Closed-Form 0.0114 $\pm$ 0.0007
MC 100 0.0115 $\pm$ 0.0006
MC 1000 0.0118 $\pm$ 0.0007
We consider Heteroscedastic Classifiers (HET) [5], Spectral-Normalized Gaussian Processes (SNGP) [20], and last-layer Laplace-Approximated Networks [8] as backbones supplying logit-space Gaussian distributions. The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [9], CIFAR-10, and CIFAR-100 [17] on five metrics aligning with practical needs from uncertainty estimates [26]. Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples. For ImageNet, we treat ImageNet-C [13] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10 and CIFAR-100, we use the CIFAR-10C corruptions.
For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [14] and Mean Field [21] approximations, as well as MC sampling, in Section 6.2.
To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones. For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [37] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al. [27]. For Vision Transformer [11] experiments, refer to Section K.2. On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The CIFAR-100 experiments train WideResNet-28-5 models for 200 epochs. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [33] in Weights & Biases [3].
The main paper focuses on ImageNet results to highlight the scalability of our framework and only shows proper scoring results for CIFAR-10. For a complete overview of CIFAR-10 and CIFAR-100 results, refer to Appendix J and Section K.3, respectively.
In all plots of the main paper and the Appendix, the error bars are calculated over five independent runs with different seeds and show the minimum and maximum metric value.
Table 2: Comparison of OOD AUROC results on ImageNet using severity level one for different uncertainty estimators on the three best-performing (method, activation) pairs: Sigmoid Laplace, NormCDF HET, and NormCDF Laplace. The best estimator is derived from the moment-matched Dirichlet distributions in all three cases. MC estimates use 1000 samples.
Method Mean $\pm$ 2 std
Sigmoid Laplace
Dirichlet Mutual Information 0.6388 $\pm$ 0.0005
Closed-form Predictive Entropy 0.6353 $\pm$ 0.0003
Monte Carlo Expected Entropy 0.6321 $\pm$ 0.0001
NormCDF HET
Dirichlet Expected Entropy 0.6353 $\pm$ 0.0021
Monte Carlo Expected Entropy 0.6281 $\pm$ 0.0020
Closed-form Predictive Entropy 0.6277 $\pm$ 0.0021
NormCDF Laplace
Dirichlet Expected Entropy 0.6321 $\pm$ 0.0009
Monte Carlo Predictive Entropy 0.6296 $\pm$ 0.0012
Closed-Form Predictive Entropy 0.6296 $\pm$ 0.0012